OpenStudy (beth12345):

A proton of mass 1.67x10^-27kg, travelling with a speed of 1.0x10^7m/s, collides with a helium nucleus at rest. The proton rebounds straight back with a speed of 6.0x10^6m/s while the helium nucleus moves forward with a speed of 4.0x10^6m/s. What was the momentum of the helium nucleus after the collision?

OpenStudy (beth12345):

\[2.67 \times 10^{-20} kg \times m/s\] is the answer I'm given, I just don't know how to get it.

OpenStudy (anonymous):

mass of helium nucleus=4 amu =4*1.67*10^-27 kg =6.68*10^-27 kg linear momentum=mass *velocity=(6.68*10^-27)*4*10^6 kg-m/sec =2.672*10^-20 kg-m/sec

OpenStudy (beth12345):

I don't know why it's multiplied by 4, and im also not sure what the amu stands for

OpenStudy (anonymous):

momentum will be conserved, which is the nature of all collisions. Therefore, \[p_p = p_p' + p_n'\] where the prime indicates the momentum after the collision. This can be expanded as\[v_p m_p = v_p' m_p + p_n'\]We can solve this for \(p_n'\) without even having to know the mass of the nucleus! Remember that since the proton rebounds, \(v_p' \lt 0\)

OpenStudy (beth12345):

ok

OpenStudy (beth12345):

thanks :)